# The Silent Majority: Demystifying Memorization Effect in the Presence of Spurious Correlations

### v2

@@ 3.2.2 Removing Spurious Memorization Improves Group Robustness
In this study, our primary objective is to investigate whether mitigating spurious memorization can lead to an enhancement in the test accuracy of minority groups. The findings are illustrated in Figure 6, where we compare the Worst Group Accuracy (WGA) between standard ERM training and our proposed framework. The WGA for ERM training is evaluated by the testing set with the best-performing checkpoint (on the validation set) among the first 100 epochs. Notably, we observe a significant increase in WGA under all scenarios. Specifically, with the ResNet-50 backbone, we observe a significant 16.87% and 10.77% improvement in WGA for the Waterbirds and CelebA datasets, respectively. Similarly, on ViT-Small model, we observe 23.83% and 9.45% improvements in WGA.
@@ 3.2.3 Ablation on Hyper-parameters
Experimental Setup Group Accuracy
ERM N/A CE 64.02% N/A
ERM+Fine-tuning (ours) 0.01% CE 71.34% 7.32%
0.01% MSE 80.89% 16.87%
0.10% CE 77.26% 13.24%
0.10% MSE 79.91% 15.89%
Ablation on Kick-in Epoch.   In our training framework introduced in Section 3.2.1, we first pretrain the target using ERM for 40 epochs and then switch to a fine-tuning stage with loss function Eq. (3.3). In other words, our framework kicks in at epoch 40. Here we test different choices of the kick-in epoch. The result is shown in Table 2. Overall, we see that epoch 30 is not effective, while the choice of 40, 50, and 60 all return meaningful returns. This reason is that the ERM training of the target model has not converged yet at epoch 30.
Experimental Setup Group Accuracy
Training Strategy Kick-in Epoch WGA Difference from ERM
ERM N/A 64.02% N/A
ERM+Fine-tuning (ours) 30 65.89% 1.87%
40 80.89% 16.87%
50 73.68% 9.66%
60 76.23% 12.21%
Ablation on Number of Fine-tuning Epochs.   We then compare the number of epochs for the fine-tuning stage. The result is shown in Table 3. Observe that, there is no difference between the result for using 20 or 30 fine-tuning epochs. This is because the best model check-point appears within 20 epochs. The result indicates that fine-tuning for more epochs is unnecessary.
Experimental Setup Group Accuracy
Training Strategy Fine-tuning Epochs WGA Difference from ERM
ERM N/A 64.02% N/A
ERM+Fine-tuning (ours) 10 54.98% -9.04%
20 80.89% 16.87%
30 80.89% 16.87%
Ablation on Data Source for Gradient Calculation.   We compare different source of gradient in the calculation of gradient-based criterion. As introduced in Section 3.2.1, the neuron gradient is computed on a selected subset of training data. By default, this subset is chosen as the worst predicted examples by the target model in terms of the CE loss. From Table 4, we observe that, calculating the gradient on the subset of worst predicted examples from the minority groups does not show any benefit. Considering the fact that using minority groups as the gradient source requires access to the group label which is sometimes unavailable, we suggest using the full training set as the gradient source.
Experimental Setup Group Accuracy
ERM N/A CE 64.02% N/A
ERM+Fine-tuning (ours) 0.01% Minority Groups 74.45% 10.43%
0.01% Full Training Set 76.64% 12.62%
0.02% Minority Groups 65.89% 1.89%
0.02% Full Training Set 78.50% 14.48%
Ablation on Loss Term Ratios.   We compare different choice of the loss term ratio $\lambda$ in Eq. (3.3). The result is shown in Table 5.
Experimental Setup Group Accuracy
Training Strategy $\lambda$ WGA Difference from ERM
ERM N/A 64.02% N/A
ERM+Fine-tuning (ours) 0.01 65.26% 1.24%
0.2 80.89% 16.87%
1 72.43% 8.41%
10 59.97% -4.05%
Ablation on Pruning Percentage.   We compare different choice of pruning percentage for both the gradient-based criterion and magnitude-based criterion. The result is shown in Table 6.
Experimental Setup Group Accuracy
ERM N/A N/A 64.02% N/A
ERM+Fine-tuning (ours) Gradient-based 0.01% 76.64% 12.62%
0.1% 77.26% 13.24%
1% 71.50% 7.48%
10% 0.04% -63.98%
ERM+Fine-tuning (ours) Magnitude-based 0.01% 80.89% 16.87%
0.1% 79.91 15.89%
1% 69.00% 4.98%
10% 73.83% 9.81%
Ablation on Combined Pruning Criteria.   We test a mixed pruning criterion which combines the gradient-based criterion with the magnitude-based one. The result is shown in Table 7.
Experimental Setup Group Accuracy
ERM N/A N/A 64.02% N/A
ERM+Fine-tuning (ours) 0.01% 0.01% 77.26% 13.24%
0.1% 0.1% 77.57% 13.55%
1% 1% 69.22% 5.20%
10% 10% 0.18% -63.84%
@@ 5.1 Experimental Setup
Datasets and Models.   In our study, we conduct experiments on two popular benchmark datasets for spurious correlation: Waterbirds (Sagawa et al., 2020a; Wah et al., 2011), and CelebA (Liu et al., 2015). We comprehensively evaluate the extent to which spurious memorization exists in the large pre-trained models (ResNet-50 (He et al., 2016) and ViT-Small (Dosovitskiy et al., 2021)) on ImageNet (Deng et al., 2009). Note that we report the average performance of 10 independent runs with different random seeds for experiments including unstructured tracing and structured tracing. In the experiments detailed in Section 3.2, we strictly adopt the standard dataset splits for both Waterbirds and CelebA, following the setting in (Idrissi et al., 2022). Our adoption of ResNet or ViT models pre-trained on ImageNet is consistent with the main literature (Kirichenko et al., 2023; Qiu et al., 2023; Yang et al., 2023). Furthermore, the high baseline accuracy achieved by pre-trained models is critical for studying memorization, which is a focal point of our study.
@@
The code that supports the findings of this study is available from the corresponding authors upon reasonable request.
@@
The Silent Majority: Demystifying Memorization Effect
in the Presence of Spurious Correlations
@@
Waterbirds Wah et al. [2011], Sagawa et al. [2020a] Input size $224\times 224$

### v3

@@ 2.2.2 Removing Spurious Memorization Improves Group Robustness
In this study, our primary objective is to investigate whether mitigating spurious memorization can lead to an enhancement in the test accuracy of minority groups. The findings are illustrated in Figure 6, where we compare the Worst Group Accuracy (WGA) between standard ERM training and our proposed framework. The WGA for ERM training is evaluated by the testing set with the best-performing checkpoint (on the validation set) among the first 100 epochs. Notably, we observe a significant increase in WGA under all scenarios. Specifically, with the ResNet-50 backbone, we observe a significant 16.87% and 10.77% improvement in WGA for the Waterbirds and CelebA datasets, respectively. Similarly, on ViT-Small model, we observe 23.83% and 9.45% improvements in WGA. Furthermore, the statistical summary is included in Table 12 (in Supplementary materials).
@@ 2.2.3 Ablation on Hyper-parameters
Ablation on Kick-in Epoch.   In our training framework introduced in Section 2.2.1, we first pretrain the target using ERM for 40 epochs and then switch to a fine-tuning stage with loss function Eq. (2.3). In other words, our framework kicks in at epoch 40. Here we test different choices of the kick-in epoch. The result is shown in Table 5 (in Supplementary materials). Overall, we see that epoch 30 is not effective, while the choice of 40, 50, and 60 all return meaningful returns. This reason is that the ERM training of the target model has not converged yet at epoch 30.
Ablation on Number of Fine-tuning Epochs.   We then compare the number of epochs for the fine-tuning stage. The result is shown in Table 6 (in Supplementary materials). Observe that, there is no difference between the result for using 20 or 30 fine-tuning epochs. This is because the best model check-point appears within 20 epochs. The result indicates that fine-tuning for more epochs is unnecessary.
Ablation on Data Source for Gradient Calculation.   We compare different source of gradient in the calculation of gradient-based criterion. As introduced in Section 2.2.1, the neuron gradient is computed on a selected subset of training data. By default, this subset is chosen as the worst predicted examples by the target model in terms of the CE loss. From Table 7 (in Supplementary materials), we observe that, calculating the gradient on the subset of worst predicted examples from the minority groups does not show any benefit. Considering the fact that using minority groups as the gradient source requires access to the group label which is sometimes unavailable, we suggest using the full training set as the gradient source.
Ablation on Loss Term Ratios.   We compare different choice of the loss term ratio $\lambda$ in Eq. (2.3). The result is shown in Table 8 (in Supplementary materials).
Ablation on Pruning Percentage.   We compare different choice of pruning percentage for both the gradient-based criterion and magnitude-based criterion. The result is shown in Table 9 (in Supplementary materials).
Ablation on Combined Pruning Criteria.   We test a mixed pruning criterion which combines the gradient-based criterion with the magnitude-based one. The result is shown in Table 10 (in Supplementary materials).
@@ 4.1 Experimental Setup
Datasets and Models.   In our study, we conduct experiments on two popular benchmark datasets for spurious correlation: Waterbirds (Sagawa et al., 2020; Wah et al., 2011), and CelebA (Liu et al., 2015). We comprehensively evaluate the extent to which spurious memorization exists in the large pre-trained models (ResNet-50 (He et al., 2016) and ViT-Small (Dosovitskiy et al., 2021)) on ImageNet (Deng et al., 2009). Note that we report the average performance of 10 independent runs with different random seeds for experiments including unstructured tracing and structured tracing. In the experiments detailed in Section 2.2, we strictly adopt the standard dataset splits for both Waterbirds and CelebA, following the setting in (Idrissi et al., 2022). Our adoption of ResNet or ViT models pre-trained on ImageNet is consistent with the main literature (Kirichenko et al., 2023; Qiu et al., 2023; Yang et al., 2023). Furthermore, the high baseline accuracy achieved by pre-trained models is critical for studying memorization, which is a focal point of our study.
@@
The code that supports the findings of this study is available at https://github.com/aarentai/Silent-Majority.
Acknowledgements
C.Y. and J.S.D. were supported by NIH grants R01CA206180 and R01HL121226. H.D. and S.J. were supported by NSF grant DMS-1912030. The authors want to thank all the anonymous reviewers and editors for their constructive comments and suggestions that substantially improved this paper. C.Y. is the corresponding author of this paper.
Authors contribution
C.Y., H.D., and Y.M. developed and implemented new machine learning methods, benchmarked machine learning models and analyzed model behavior, all under the guidance of J.S.S., S.J., J.S.D., and performed a validation study to evaluate its effects. All authors discussed the results and contributed to the final manuscript. C.Y., H.D., and Y.M. designed the study.
Competing interests
The authors declare no competing interests.
@@ Acknowledgements
Uncovering Memorization Effect in the Presence of Spurious Correlations
@@
Figure 8: Distribution of Critical Neurons in ResNet-50 via Unstructured Tracing. This analysis utilizes a gradient-based criterion to identify critical neurons, with the color bar indicating the neuron count.
@@
Waterbirds Wah et al. [2011], Sagawa et al. [2020] Input size $224\times 224$
@@
Experimental Setup Group Accuracy
ERM N/A CE 64.02% N/A
ERM+Fine-tuning (ours) 0.01% CE 71.34% 7.32%
0.01% MSE 80.89% 16.87%
0.10% CE 77.26% 13.24%
0.10% MSE 79.91% 15.89%
Experimental Setup Group Accuracy
Training Strategy Kick-in Epoch WGA Difference from ERM
ERM N/A 64.02% N/A
ERM+Fine-tuning (ours) 30 65.89% 1.87%
40 80.89% 16.87%
50 73.68% 9.66%
60 76.23% 12.21%
Experimental Setup Group Accuracy
Training Strategy Fine-tuning Epochs WGA Difference from ERM
ERM N/A 64.02% N/A
ERM+Fine-tuning (ours) 10 54.98% -9.04%
20 80.89% 16.87%
30 80.89% 16.87%
Experimental Setup Group Accuracy
ERM N/A CE 64.02% N/A
ERM+Fine-tuning (ours) 0.01% Minority Groups 74.45% 10.43%
0.01% Full Training Set 76.64% 12.62%
0.02% Minority Groups 65.89% 1.89%
0.02% Full Training Set 78.50% 14.48%
Experimental Setup Group Accuracy
Training Strategy $\lambda$ WGA Difference from ERM
ERM N/A 64.02% N/A
ERM+Fine-tuning (ours) 0.01 65.26% 1.24%
0.2 80.89% 16.87%
1 72.43% 8.41%
10 59.97% -4.05%
Experimental Setup Group Accuracy
ERM N/A N/A 64.02% N/A
ERM+Fine-tuning (ours) Gradient-based 0.01% 76.64% 12.62%
0.1% 77.26% 13.24%
1% 71.50% 7.48%
10% 0.04% -63.98%
ERM+Fine-tuning (ours) Magnitude-based 0.01% 80.89% 16.87%
0.1% 79.91 15.89%
1% 69.00% 4.98%
10% 73.83% 9.81%
Experimental Setup Group Accuracy
ERM N/A N/A 64.02% N/A
ERM+Fine-tuning (ours) 0.01% 0.01% 77.26% 13.24%
0.1% 0.1% 77.57% 13.55%
1% 1% 69.22% 5.20%
10% 10% 0.18% -63.84%
Table 11: Unstructured Tracing on CelebA. We conduct the additional unstructured tracing experiment on CelebA. The table presents the changes in training accuracy for each group. Note that for the CelebA dataset, the minority group is denoted as $\mathcal{G}_{3}^{*}$, with the star superscript indicating its minority status.
Experimental Setup Group Accuracy % Change
Modification top-$k$ Std. Dev. $\mathcal{G}_{0}$ $\mathcal{G}_{1}$ $\mathcal{G}_{2}$ $\mathcal{G}_{3}^{*}$
zero-out 1 n/a 2.64 0.26 4.85 10.24
zero-out 2 n/a 2.63 0.23 4.77 9.23
zero-out 3 n/a 3.34 0.26 5.29 9.66
random init 1 0.005 2.68 0.26 4.86 10.34
random init 2 0.005 2.59 0.23 4.75 9.16
random init 3 0.005 3.39 0.28 5.32 9.93
random init 1 0.01 2.81 0.28 4.94 10.86
random init 2 0.01 2.87 0.26 4.96 9.98
random init 3 0.01 3.30 0.25 5.17 9.21
random init 1 0.02 2.88 0.30 4.97 11.13
random init 2 0.02 2.54 0.23 4.51 8.53
random init 3 0.02 3.57 0.31 5.37 10.82
Table 12: Summary of Experiments in Sec. 2.2.2. This table presents the mean and standard deviation of Worst Group Accuracy (WGA) for our proposed method, computed over 10 random seeds. The mean values reported here are consistent with those shown in Figure 6.
WGA (%)
Setting ERM Gradient Magnitude
ResNet on Waterbirds 64.0$\pm$0.45 77.3$\pm$4.91 80.9$\pm$4.36
ResNet on CelebA 47.0$\pm$1.25 57.8$\pm$3.58 55.6$\pm$2.01
ViT on Waterbirds 52.7$\pm$2.79 69.6$\pm$5.52 76.5$\pm$5.74
ViT on CelebA 47.2$\pm$1.59 56.7$\pm$3.74 52.8$\pm$1.32
Table 13: Comparison between Changes in Training Accuracy and Changes in Test Accuracy. We compare the percentage change in the training accuracy and test accuracy on Waterbirds. The results show that for the minority groups $\mathcal{G}_{1}^{*}$ and $\mathcal{G}_{2}^{*}$ under almost all experimental setups, the test accuracy is affected significantly less than the training accuracy when modifying the critical neurons. Specifically, the test accuracy for minority groups remains quite stable after modifying critical neurons, as opposed to the training accuracy. This result provides additional evidence for the memorization role of the critical neurons. Note that for the Waterbirds dataset, the minority group is denoted as $\mathcal{G}_{1}^{*}$ and $\mathcal{G}_{2}^{*}$, with the star superscript indicating their minority status.
Experimental Setup Group Accuracy % Change (training set) Group Accuracy % Change (test set)
Modification Top-$k$ Std. Dev. $\mathcal{G}_{0}$ $\mathcal{G}_{1}^{*}$ $\mathcal{G}_{2}^{*}$ $\mathcal{G}_{3}$ $\mathcal{G}_{0}$ $\mathcal{G}_{1}^{*}$ $\mathcal{G}_{2}^{*}$ $\mathcal{G}_{3}$
zero-out 1 N/A 1.00 4.90 3.57 0.38 0.93 0.89 1.25 0.16
zero-out 2 N/A 0.57 3.07 7.15 0.18 1.51 2.48 0.31 0.62
zero-out 3 N/A 0.94 1.63 10.72 0.28 1.77 2.17 0.31 0.16
random noise 1 0.005 0.08 0.34 0.67 0.01 0.07 0.31 0.12 0.03
random noise 2 0.005 0.08 0.78 0.89 0.03 0.09 0.23 0.22 0.05
random noise 3 0.005 0.06 0.76 0.89 0.03 0.08 0.35 0.22 0.07
random noise 1 0.01 0.06 0.61 0.89 0.02 0.07 0.39 0.29 0.02
random noise 2 0.01 0.07 0.60 0.20 0.02 0.08 0.40 0.34 0.12
random noise 3 0.01 0.03 1.52 0.71 0.03 0.16 0.27 0.22 0.24
random noise 1 0.02 0.09 1.29 0.67 0.06 0.13 0.54 0.16 0.16
random noise 2 0.02 0.11 1.44 1.59 0.01 0.17 0.49 0.22 0.28
random noise 3 0.02 0.09 1.84 2.32 0.06 0.31 0.47 0.29 0.33
Table 14: Raw Values of Change in the Group Accuracy. This table provides the raw values for the changes in group accuracy corresponding to the experimental results in Figures 4 and 5. It clearly demonstrates that for minority group $G_{1}^{*}$ and $G_{2}^{*}$ training accuracy consistently decreases across all experimental settings and for every choice of $k$ in top-$k$. Note that for the Waterbirds dataset, the minority group is denoted as $\mathcal{G}_{1}^{*}$ and $\mathcal{G}_{2}^{*}$.
Group Accuracy % Change
top-$k$ $\mathcal{G}_{0}$ $\mathcal{G}_{1}^{*}$ $\mathcal{G}_{2}^{*}$ $\mathcal{G}_{3}$
Figure 4 (left) top1 1.25 -8.92 -6.08 -0.18
top2 1.16 -11.80 -6.97 1.83
top3 -0.36 -6.80 -3.04 1.71
Figure 4 (right) top1 -1.30 -6.95 -2.19 -0.18
top2 0.90 -2.84 -3.38 0.10
top3 -2.63 -10.03 -3.54 -0.37
Figure 5 (left) top1 -0.07 -0.30 -0.6 -0.01
top2 -0.08 -0.78 -0.89 -0.03
top3 -0.07 -0.84 -0.89 -0.04
Figure 5 (right) top1 -0.86 -3.20 -3.97 0.30
top2 -0.59 -0.18 -6.75 -0.17
top3 -0.89 -1.09 -9.13 -0.28